Can you tell me how to set this problem up?
1. Suppose the club earned a total of $181 for washing 49 vehicles. If they charge $3 for a car and $5 for a truck, how many of each vehicle were sold?

Found 2 solutions by priya, stanbon:
Answer by priya(51)

DO U WANT TO KNOW THE NUMBER OF EACH VEHICLE WASHED R SOLD???
AND DO U MEAN THAT THE MONEY EARNED WAS WASHING VEHICLES(BOTH CARS AND TRUCKS R CARS ALONE)?????
CHECK UR QUESTION PLZ.

Answer by stanbon(48568)

Let number of cars washed be "c"
Money from that work is "3c"
Let number of trucks washed be "t"
Money from that work is "5t"
EQUATION9:
1) c + t = 49
2) 3c+5t=181
Multiply thru 1t equation by 3 to get:
1* 3c+3t= 147
Subtract 1* from 2) to get:
2t=34
t=17 (# of trucks washed)
c=49-17= 32 (# of cars washed)
Cheers,
Stan H.